what are the total formulaes of atomic structure in bohr atumicc model

Grade:12

2 Answers

Umakant biswal
7 years ago
@ harshith 
your query are not clear , i guess u want to know about the formula related to the bohrs model of atom , 
if i am correct then below are the explanation and i am enlisting all the formula here , 
lambda = h/mv (calculation of wavelength ) 
2 * pi * r = n lambda 
L= mvr = hr/ lambda = nh / 2 * pie 
U= – ZE SQUARE / 4 PIE EFSILON NOT R ( POTENTIAL ENERGY ) 
T+U = – ze square / 8 * pie * efsilon not * r 
mv square / 2=  n^2 h ^2 / 8 pie ^2 mr^2 = ze^2 / 8 pie efsilonnot r 
r = n^2 h^2 efsilon not / z pie me^2 ( bohr radius ) 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
Mudit
7 years ago
mvr=nh/2pi
Kq1q2/r2=mv2/r
Energy of electron=-13.6*Z2/n2 eV
Radius=0.529*n2/z A
Velocity of electron=2.18*106Z/n
Rvolutions per sec =v/2r(pi)
Time for one revolution=2r(pi)/v
totalenergy of electron=-KE=PE/2
plz approve if u find it helpful.